OpenStudy (anonymous):

Solve: 4y − y + 11 = 38 9 −9 3.8 8.6

OpenStudy (anonymous):

3y = 27 y = 9

OpenStudy (diyadiya):

4y − y + 11 = 38 3y+11=38 Subtract 11 from both sides 3y +11-11=38-11 3y=27 divide both sides by 3 y =?

OpenStudy (anonymous):

9.

OpenStudy (anonymous):

right...